OpenStudy (anonymous):

The velocity of sound in air is given by the equation where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the temperature when the velocity of sound in air is 395 meters per second. Round to the nearest degree. A. 7,528ºC B. 517ºC C. 117ºC D. 8,074ºC

hartnn (hartnn):

which equation ? you just to put v= 395 and find 't' from that equation.

OpenStudy (anonymous):

thats all it says

hartnn (hartnn):

you must be having an equation with you...

hartnn (hartnn):

ok, then you may use this equation taken from wikipedia, v= 331.3 - 0.606 t just plug in v = 395 and find t=...?

OpenStudy (anonymous):

thats not one of the answers

hartnn (hartnn):

but its nearer to one...

OpenStudy (anonymous):

let v=speed of sound, t=temperature in Celcious scale. v=331*sqrt(1+t/273.15) => t=273.15*[(v/331)^2-1] put v=395, t=115.8408..................... so nearest value is answer C